The pH of a solution is 3.72. What is the OH⁻ concentration in the solution?


1
Expert's answer
2022-02-14T21:56:02-0500

pH= 14-pOH

"pH= 14-pOH"

But pOH= 14- 3.72= 10.28

[OH-]= 10-poH

= 10-10.28